A fisherman uses a fishing rod to pull a fish of weight 6.5 kg out of a river. The fish is then suspended vertically on the end of a fishing line of uniform circular cross section. The line has a breaking stress of 65 MPa.
Calculate:
(a)
The minimum diameter of the fishing line required if the line does not break.
(b)
The extension per metre of this line under these conditions.
1
Expert's answer
2020-07-02T17:15:11-0400


(a)


"\sigma=\frac{F}{S}=\frac{4F}{\pi d^2}\to d=\sqrt{\frac{4F}{\pi\sigma}}=\sqrt{\frac{4\cdot 6.5\cdot9.81}{3.14\cdot 65\cdot10^6}}=0.0011m"


(b)


"\sigma=E\frac{\Delta l}{l}" if "l=1 m" then "\Delta l=\frac{\sigma}{E}"


Assume that "E=1.4\cdot10^9Pa"


So, we get


"\Delta l=\frac{\sigma}{E}=\frac{65\cdot10^6}{1.4\cdot10^9}=0.046 m"